Find the dy/dx of each of the following:

1.y=〖5x〗^3-〖4x〗^2+3x-6


2.y=∛x+4/x+√x


3.y=(3-x^2 )(x^3-x+1)


4.y=(x^2+1)(x+5+1/x)


5.y=(5x+1)/(2√x)


1
Expert's answer
2022-03-14T04:32:14-0400

1.dy/dx=375x2-32x+3

2."dy/dx=\frac{1}{3\sqrt[3]{x^2}}-\frac{4}{x^2}+\frac{1}{2\sqrt{x}}"

3."y=3x^3-3x+3-x^5+x^4-x^2"

"dy\/dx=9x^2-3-5x^4+4x^3-2x"

4."y=x^3+5x^2+x+x+5+1\/x"

"dy/dx=3x^2+10x+2-\frac{1}{x^2}"

5."dy/dx=\frac{5}{4\sqrt{x}}-\frac{1}{4\sqrt[3]{x^2}}"